1) Interest rates are __________ to the level of loanable funds available to borrowers. Select one:...

1)

Interest rates are __________ to the level of loanable funds available to borrowers. Select one: A. perfectly correlated B. inversely related C. totally disproportional D. completely unrelated

2) If inflation is anticipated to be 3.00% per year and the nominal interest rate is currently 6.50%, which of the following represents the real rate of interest under conditions of rational expectations? Select one: A. 9.50% B. 3.50% C. 6.50% D. 3.00% E. None of the above
Answer #1

1.

Option A is correct

Perfectly correlated.

Explanation:

Under simplified model of microeconomics, as the interest rates increase more and more lenders will give loans to earn higher interest rate on their funds. But at a certain level, number of borrowers seeking loan will decrease due to higher interest rate.

2.

As per Fischer equation,

(1 + N) = (1 + R) × (1 + I)

As an approximation to Fischer Equation

N = R + I

R = N - I

R = 6.50 - 3.00

Real Interest Rate = 3.50%

Option B is correct.